OpenStudy (anonymous):

(5w + 2) (8w +5) A: 40w^2 + 41w - 10 B: 40w^2 + 41w + 10 C: 40w^2 - 41w - 10 D: 40w^2 - 41w +10

OpenStudy (texaschic101):

are you familiar with the FOIL method ?

OpenStudy (azureilai):

You unfoil the equation. 5w x 8w 5w x 5 8w x 2 2 x 5 Multiply all the terms, then simplify and add like terms

OpenStudy (texaschic101):

FOIL -- first, outside, inside, last F -- multiply first terms in each set 5w * 8w = ? O -- multiply outside terms in each set 5w * 5 = ? I -- multiply inside terms in each set 2 * 8w = ? L -- multiply last terms in each set 2 * 5 = ? put it all together, then collect like terms and you will have your answer

OpenStudy (texaschic101):

so what 5w * 8w ??

OpenStudy (anonymous):

Okay I understand, so the answer would be B right?

OpenStudy (texaschic101):

40w^2 + 25w + 16w + 10 collect like terms 40w^2 + 41w + 10 yes...answer is B

OpenStudy (anonymous):

Thank you for your help

OpenStudy (anonymous):

Okay I understand, so the answer would be B right?